OpenStudy (anonymous):

what is the integration of dy/dx=sec^2y

OpenStudy (fortytherapper):

So, is iy \[\int\limits_{}^{}\sec^2ydy\]

OpenStudy (fortytherapper):

it*

OpenStudy (anonymous):

it is not like this

OpenStudy (anonymous):

x=1/2y+1/4sin2y

OpenStudy (anonymous):

rahman hope you are ok with the answer

OpenStudy (irishboy123):

yep, agreed @Gyimah can't see us getting that into a form that says \(y(x) = f(x)\) !! so you can go \(x' = \cos^2 y \implies \int dx = \int \dfrac{\cos 2y + 1}{2} \; dy\) which will give a form that says \(x(y) = f(y)\) but don't forget the integration constant :-)

OpenStudy (anonymous):

thanks